solve the following equation for x.
2/5=x-2/20

Answer by checkley71(8403)

2/5=x-2/20 or 2/5=x-1/10 or 2/5+1/10=x or 4/10+1/10=x or 5/10=x or x=1/2
proof 2/5=1/2-2/20 or 2/5-1/2+2/20=0 or 2/5-1/2+1/10=0 or 4/10-5/10+1/10=0 or
(4-5)/10+1/10=0 or -1/10+1/10=0 or 0=0